OpenStudy (anonymous):

Suppose that waiting times in a medical office are continuously and uniformly distributed between 1 and 13 minutes. (a) what is the probability of waiting exactly 9 minutes? (b) what is the standard deviation of the waiting time?

OpenStudy (anonymous):

i guess the first one is zero, since it is a continuous distribution

OpenStudy (anonymous):

standard deviation is square root of variance, which if i remember correctly is \(\frac{b-a}{\sqrt{12}}\)

OpenStudy (anonymous):

in your case \[\frac{13-1}{\sqrt{12}}=\frac{12}{\sqrt{12}}=\sqrt{12}\]